Find the variance of the following data. Round your answer to one decimal place.

x7891011
P(X=x)0.10.20.20.30.2
Answer #1

Solution

x P(x) x * P(x) x2 * P(x)
7 0.1 0.7 4.9
8 0.2 1.6 12.8
9 0.2 1.8 16.2
10 0.3 3 30
11 0.2 2.2 24.2
Sum 1 9.3 88.1

Mean = \sum X * P(X)

\mu = 9.3

Variance =

= \sum X 2 * P(X) - \mu 2

= 88.1 - (9.3)2 = 1.6